Operation Summary for November 20, 2009 - November 23, 2009 |
The Integrated Luminosity for the period from 11/16/09 to 11/23/09 was 330.16 pb-1. Friday Nov. 20 The day shift began with Operations monitoring the Tevatron (TeV) store 7370 (37.87E30) and the Recycler stash (292.6E10), with the Antiproton Source stacking (13.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7370 at 8:33 AM. D0 personnel accessed their collision hall. A Linac RF station (LRF4) tripped off just prior to 9:30 AM; an investigation of problems with the crowbar circuit began and continued throughout the shift. Preaccelerator personnel successfully pumped down the I- Source vacuum. Booster personnel repaired BRF9. During the evening shift, Linac personnel completed their LRF4 repairs and beam was re-established by 5:45 PM. Operators checked the TeV orbits and then put the TeV into HEP shot setup at 8 PM. Operations established store 7373 at 9:45 PM with an average initial luminosity (AIL) of 257E30. Instrumentation indicated that the Pbar target stopped rotating towards the end of the shift. Operations halted stacking and contacted Pbar personnel. Saturday Nov. 21 The midnight shift began with Operations monitoring store 7373 (168.45E30), the stash (92.5E10), and the stack (11.3E10), with Pbar not stacking due to target problems, and with NuMI and MiniBooNE taking beam. During the midnight shift, Pbar personnel investigated the Pbar target rotation problem and determined that the target was indeed rotating, but the readback was in error. Target personnel provided an alternative method of verifying target rotation for the time being. The MI-40 abort kicker tripped off and would not reset. Operations called in EE Support personnel to investigate. The Pelletron tripped off a few times; Operations contacted Recycler personnel for guidance. During the day shift, EE Support personnel replaced the MI-40 kicker thyratron tube and Operations re-established beam to the Pbar Source and NuMI by 2 PM. Store 7373 aborted when the controls crate at TeV sector B1 failed, which caused the vacuum valves to close. Operators replaced a fuse in the crate. Operations put the TeV into shot setup at 3:45 PM. During the evening shift, The TeV shot setup ended at 5 PM with a quench at TeV sector A4 during the low beta magnet squeeze. The Pelletron had developed a serious high-voltage regulation problem during shot setup. Operations called in Electron Cooling personnel to transfer the SF6 gas out of the Pelletron; when completed they will access and replace the triode. Operators had the TeV checked out and turn-on underway by 7:41 PM. Operations put the TeV through a dry squeeze at 11:47 PM. Sunday Nov. 22 The midnight shift began with the TeV in a dry squeeze, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operators put the TeV through a wet squeeze and then checked the TeV orbits. Operations put the TeV into shot setup at 1:15 AM. Operations established store 7380 at 2:45 AM with AIL of 90E30. The Pelletron gas removal continued. The number of stacking pulses was drastically reduced due to the lack of cooling in the Recycler and the resulting 150E10 stash limit. During the day shift, Pelletron gas removal continued. The H- Source vacuum valve (between the turbo and ion pump) was opened. Operations terminated store 7380 at 4 PM. During the evening shift, Operations put the TeV into shot setup at 4:15 PM, this was after a short delay to tighten the belts on the TeV sector F4 wet engine. Shot setup then came to an abrupt halt due to CDF crate problems. Operations put the TeV into a dry squeeze until CDF recovered their calorimeter system readbacks. Operations put the TeV into shot setup again at 7:20 PM. Operations established store 7382 at 8:45 PM with an AIL of 92E30. Operations reported that the Pelletron gas transfer was completed by 11:16 PM; a Pelletron person had the high-voltage regulation repairs completed by 11:57 PM. Monday Nov. 23 The midnight shift began with Operations monitoring store 7382 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, the Pelletron resumed normal operations by 2:30 AM. Operations terminated store 7382 at 3:45 AM. Operation put the TeV into shot setup and established store 7383 at 5:20 AM with an AIL of 194E30. The Plans for Monday The plans for today are to stack, stash, and store. Switchyard personnel will send beam to the Switchyard absorber and then on to the Meson MTest area. |
|
More Information Comments and Suggestions |
| last modified 11/23/2009 email Fermilab |
| Security, Privacy, Legal | |